OpenStudy (inexilewetrust):

2(length)(width)+2(length)(height)+2(width)(height) Length=4 meters and height=3 meters If the area is 164 meters, what is the width?

OpenStudy (3mar):

May I help?

OpenStudy (ben00):

do 164 divided by (4 x 3) to get the answer.

OpenStudy (inexilewetrust):

I did, that isn't the right answer.

OpenStudy (inexilewetrust):

2(4)(width)+2(4)(3)+2(width)(3)=164.

OpenStudy (inexilewetrust):

That is as far as I got in getting the help I need.

OpenStudy (3mar):

Can I?

OpenStudy (inexilewetrust):

@3mar sure, this is open to anyone.

OpenStudy (3mar):

I know but I need your attention!

OpenStudy (inexilewetrust):

Oh course. I am all eyes and ears.

OpenStudy (3mar):

"2(4)(width)+2(4)(3)+2(width)(3)=164." this is what you got, right?

OpenStudy (inexilewetrust):

Yes.

OpenStudy (3mar):

Good! Next: more simplification would be carried out: 2(4)(width)+2(4)(3)+2(width)(3)=164. 2*4=8 8(width)+2(4)(3)+2(width)(3)=164. 2*3*4=24 8(width)+24+2(width)(3)=164. 2*3=6 8(width)+24+6(width)=164. then add similar terms together: 8 width + 6 width = 14 width 14(width)+24=164. 164-24=140 14(width)=140 width=140/14=10 Done! Hope that helps

OpenStudy (inexilewetrust):

Why did you multiply by 4?

OpenStudy (3mar):

where?

OpenStudy (inexilewetrust):

2*3*4=24

OpenStudy (3mar):

we ahve two faces with dimensions for each 3 and 4 What you think their areas will be? Look at this. It would show more that writing!

OpenStudy (inexilewetrust):

Okay, thanks!

OpenStudy (3mar):

You are welcome! Any Help... Anytime.

OpenStudy (3mar):

Thank you for the medal! Did you get what you are looking for?

OpenStudy (inexilewetrust):

Yup! I already knew the answer, I just didn't know the steps I needed to take to get there.

OpenStudy (3mar):

Did you now?

OpenStudy (inexilewetrust):

No. But thanks to you, I do now.

OpenStudy (ben00):

I deserved the medal!